OpenStudy (anonymous):

Solve triangle ABC given that A = 56 degrees, B = 57 degrees, and b = 9.

hero (hero):

Use the fact that the interior angles of a triangle add to 180 degrees to find Angle C. Use law of sines to find the the lengths of sides a and c: \[\frac{\sin A}{a} = \frac{\sin B}{b} = \frac{\sin C}{c}\]

OpenStudy (anonymous):

So: sin 56/a = sin 57/9 = sin 67/c right?

hero (hero):

Yes, that appears to be correct. Now to solve for a and c, simply pair the following equations together: \[\frac{\sin 56}{a} = \frac{\sin 57}{9}\] \[\frac{\sin 57}{9} = \frac{\sin 67}{c}\]

hero (hero):

Make sure you are in degree mode if you use a calculator to solve.

OpenStudy (anonymous):

What do you mean degree mode?

OpenStudy (anonymous):

on your calculator there is a button that says Radians and if you click on it will say degree's where it said Radians

OpenStudy (anonymous):

I'm not finding a Radians button :/

hero (hero):

What kind of calculator do you have?

OpenStudy (anonymous):

TI 83 Plus

hero (hero):

Do you see a mode button?

OpenStudy (anonymous):

yes!

hero (hero):

Go to it and tell me what you see after clicking it.

OpenStudy (anonymous):

I see the radian / degree option and it's already on degree

OpenStudy (anonymous):

thanks so much for the help!

hero (hero):

You're welcome